HOW CAN YOU MAKE A 10% MIXTURE FROM 5% AND 50% MIXTURES
Answer by mananth(16946)

let 5% be x
let 50% be y
x+y=1
0.05x+0.5y=0.1
multiply by 100
5x+50y=10
..
5x+5y-5x-50y=5-10
-45y=-5
y=1/9
..
x=1-1/9
x=8/9
..
for every part of 50% acid you have to add 8 parts of 10% acid